A zero-coupon bond that will pay out $1,000 ten years from today sells at a price of $500. What is the Yield to Maturity on the bond? (Express your answer as the number that goes before the percent sign and use two decimal places of precision, e.g, if the yield to maturity is 18.415%, write 18.42)
Answer #1
Bond's price 500
Redemption price 1000
The time period in years 10
Yield R
Redemption price= Bond price * (1+R)^10
1000= 500* (1+R)^10
(1+R)^10= 1000/500
(1+R)^10= 2
(1+R)= 2^(1/10)
(1+R)= 1.071773
R= 0.071773
R= 7.18